OpenStudy (anonymous):

What is the rule for 1,2,4,7,11

OpenStudy (anonymous):

Everytime you move to the next number you add 1, then 2, then 3, etc ex 1+1=2; 2+2=4 4+3=7 7+4=11 11+5=16 16+6=22 22+7=29 29+8=37